Find the sum of n terms of the series
1+2+4+7+11+16++22+...+(nth term)

Answer by Edwin McCravy(20056)

Find the sum of n terms of the series
1+2+4+7+11+16++22+...(nth term)

 
First find the successive sums of the 7 terms given:
 
1 = 1 = S(1)
1 + 2 = 3 = S(2)
1 + 2 + 4 = 7 = S(3)
1 + 2 + 4 + 7 = 14 = S(4)
1 + 2 + 4 + 7 + 11 = 25 = S(5)
1 + 2 + 4 + 7 + 11 + 16 = 41 = S(6)
1 + 2 + 4 + 7 + 11 + 16 + 22 = 63 = S(7)
 
 
Write those in a list beside the 1st 7 positive 
integers:
 
n  S(n)
1   1
2   3
3   7
4  14
5  25
6  41
7  63
 
Form the column of 1st differences by
writing beside each number the difference 
between it and the number just below it 
 
n  S(n) d1 
1   1   2
2   3   4
3   7   7
4  14  11
5  25  16
6  41  22
7  63
 
Form the column of 2nd differences by
writing beside each number the difference between
it and the number just below it 
 
n  S(n) d1 d2  
1   1   2   2  
2   3   4   3  
3   7   7   4  
4  14  11   5 
5  25  16   6
6  41  22
7  63
 
Form the column of 3rd differences the
same way.  
 
n  S(n) d1 d2 d3  
1   1   2   2  1
2   3   4   3  1
3   7   7   4  1
4  14  11   5  1
5  25  16   6
6  41  22
7  63
 
Since the column of 3rd differences is
constant, we assume a 3rd degree equation
for S(n):
 
S(n) = An² + Bn² + Cn + D
 
Since we have only 4 unknowns, we only need
the first four sums:

S(1) =  1 = A*1³ + B*1² + C*1² + D
S(2) =  3 = A*2³ + B*2² + C*2² + D
S(3) =  7 = A*3³ + B*3² + C*3² + D
S(4) = 14 = A*4³ + B*4² + C*4² + D
 
Simplifying, we have this system of 
4 equations in 4 unknowns:
 
  A +   B +  C + D =  1
 8A +  4B + 2C + D =  3  
27A +  9B + 3C + D =  7
64A + 16B + 4C + D = 14
 
We solve this system of equations and get
 
A = 1/6, B = 0, C = 5/6, D = 0
 
So the equation

S(n) = An² + Bn² + Cn + D

becomes

S(n) = (1/6)n³ + (5/6)n
 
or simplified in factored form:
 
S(n) = n(n²+5)/6
 
Edwin